

********************************************************************
UARS WEEKLY STATUS REPORT - 18 MAR 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 51% of
each day's data amounting to nearly 472 minutes of solar observing time
per day.
***********
SUSIM
***********
During this past week, SUSIM successfully gathered daily mid (1.1 nm)
resolution, MgII, occultation, solar pointing offset, and continuous
monitoring scans for time periods with UARS telemetry or tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 March 2005 (Orbit 73782, GMT Day 064) through
11 March 2005 (Orbit 73887, GMT Day 070).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 067/2318 73842 AXIS 1 & 2 OFF
PEM 068/1341 73851 ZEPS/NEPS DAY-ONLY
HALOE 068/1400 73851 POWER ON
HRDI 070/1833 73884 TO LOW POWER MODE
The beta angle ranged between the angles of 61.4 degrees to 43.6
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 3.0 degrees on 21 March 2005 (DOY 080).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 4.35 to 3.97 degrees C. Additional
battery performance data for each battery on 04 March 2005 and 11
March 2005 are provided in Appendix D.
The clock error ranged from 10.9 to -10.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 02
HALOE = 01
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. The MPS
continues to utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. All
test results were positive and the official release of the new
software was installed on the backup CMS. With the install complete,
the FOT will run a few simulated days of operations to confirm the
software. Phase 1 of the simulations has been completed. Some
additional testing/verification will be required next week, as well as
uplink tests to the spacecraft. If all goes well, then the new
release will be put onto the operational CMS.
MEETINGS
The FOT attended the biweekly UARS clock rollover meeting on 01 March
2005 to the timeline of activities and current action items to
complete this task.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 March 2005 (DOY 080). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 March 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73784 064/0328 05-060 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73830 067/0503 05-061 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
73842 067/2318 05-062 PEM AXIS 1&2 OFF
73851 068/1341 05-063 PEM ZEPS/NEPS DAY-ONLY
73851 068/1400 05-064 HALOE POWER ON
73875 070/0328 05-065 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73884 070/1833 05-066 HRDI TO LOW POWER MODE
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 March 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 March 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73784 064/0354 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73830 067/0503 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
73875 070/0333 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 March 2005
04 March 2005 - GMT Day 063 Beta = 63.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 201.6/ 89.6 +9.06/+7.74 +20.8/-14.4 27.2 1.10 16.8
3 -22.4/-100.8 +5.15/+4.51 +18.8/-13.2 27.0 1.06 15.9
11 March 2005 - GMT Day 070 Beta = 43.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------ -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.80/-2.74 +00.0/+00.0 07.0 0.00 0.0
2 218.4/ 78.4 +11.76/+10.40 +26.8/-15.2 26.7 1.10 19.7
3 -33.6/-134.4 +7.41/ +6.43 +24.0/-13.2 26.7 1.06 18.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 March 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT 73836 067/1437 30832 Generic Late Acquisition
05-013 FOT 73876 070/0549 38897 Drop Out on the Return
Link
05-014 FOT 73881 070/1351 30824 Data Dropout
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-011 FOT 73699 058/1213 38687 Negative Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT See List Mar 05 30832 Generic Late Acquisition
05-013 FOT 73876 070/0549 38897 Drop Out on the Return
Link
05-014 FOT 73881 070/1351 30824 Data Dropout